OpenStudy (anonymous):

Which order pairs are the solution to the problem? A. (10,6) and (1,5) B. (10,4) and (5,1) C. (10,6) and (5,1) D. (10,5) and (5,1)

OpenStudy (anonymous):

@badmood @johnweldon1993 @HackberryAbby18

OpenStudy (anonymous):

@CallMeKiki @AQ99 @brianna98bri @Ciarán95 @dejmccray @EclipsedStar @Firejay5 @jagr2713

OpenStudy (dejmccray):

what problem?

OpenStudy (anonymous):

y=x-4 sorry I forgot to put that

OpenStudy (dejmccray):

its fine plug order pairs in X,y 10-4=?

OpenStudy (anonymous):

10-4=5

OpenStudy (dejmccray):

10-4 is?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

The answer would be 5 for the question you just asked

OpenStudy (dejmccray):

Close it's actually 6 close so the order pair (10,6) is correct cause 10-4=6 which correct now next (1,5) 1-4=-3 which is not 5 so 1,5 is not correct since we know (10,6) is correct what other choice have the order pair (10,6)?

OpenStudy (anonymous):

@dejmccray yeah I had meant to say 6 not 5.

OpenStudy (dejmccray):

its fine when you see what other choices have (10,6) you see the answer is C but ordered pairs fit c;